# MULTI-SPECTRAL REMOTE SENSING IMAGE RETRIEVAL USING GEOSPATIAL FOUNDATION MODELS

### v1

@@ 1 Introduction
Geospatial Foundation Models (GeoFM), for instance Prithvi [9], open new possibilities because they are pre-trained on a vast amount of multi-spectral data. The models have been successfully applied to various tasks, including flood and wildfire segmentation [9]. GeoFMs can utilize multi-spectral data for CBIR without requiring fine-tuning.
@@ 2 Related work
Image retrival in remote sensing has evolved significantly, shifting focus from traditional metadata-based methods to more advanced techniques that use computer vision to analyze the image content [4]. Among these developments, deep hash networks have played a central role [5, 8, 10, 11]. These models compress images into hash codes which are stored in a database and used for calculating distances. Researchers explored various learning techniques, such as contrastive learning [8] and metric learning [12], with the aim of minimizing the need for extensive training data.
@@
GeoFMs, such as Prithvi and Presto [18], are characterized by their ability to process multi-spectral data. Their pre-training on diverse satellite images overcomes the limitations of general models and benefits earth observation tasks [9].
@@ 3 Approach
The CBIR task requires identifying and retrieving images from a large database based on their visual similarity to a given query image. This process requires an efficient mechanism to compare and rank the database images in terms of their similarity to the query.
@@
Model Method
BigEarthNet-43
BigEarthNet-19
ForestNet-12
ForestNet-4
Mean
@@
Binary emb. 97.44 97.83 43.28 59.85 74.6
Trivial hash
82.75
84.26
34.91
51.71
63.41
LSH
68.36 ± 4.96
71.18 ± 4.66
28.98 ± 1.85
45.63 ± 1.39
53.47
Prithvi-100M-RGB Embedding
92.15
93.17
38.65
53.85
69.46
Binary emb.
91.38
92.43
38.11
53.31
68.81
Trivial hash
73.36
75.47
32.35
50.47
57.91
LSH
53.09 ± 8.43
55.21 ± 8.61
29.02 ± 1.45
45.42 ± 1.26
45.65
ViT-B/16-RGB Embedding
89.31
90.21
38.92
56.49
68.73
Binary emb.
88.71
89.7
39.19
57.01
68.65
Trivial hash
75.74
77.34
33.01
49.95
59.01
LSH
76.52 ± 0.95
78.02 ± 0.92
32.77 ± 1.32
49.13 ± 2.09
59.08
Table 1: mAP@20 results for all models and datasets. We highlight the best-performing method in bold and underline the second-best one. LSH is reported with a 95% confidence interval based on five seeds as this method uses random planes.
@@ 4.1 Models
The underlying GeoFM is Prithvi-100M with a model input of 224x224 pixels. The model is a Vision Transformer (ViT) [21] with 100M parameters. Prithvi processes six bands, unlike the three RGB channels of the vanilla ViT. The pre-training includes a subset of the Harmonized Landsat-Sentinel (HLS) dataset consisting of images from the Landsat 8 and Sentinel-2 satellites. We also report the results of Prithvi only processing the RGB channels (Prithvi-100M-RGB). The model input of the infrared channels is set to zero. For comparison, we include the vanilla ViT-B/16 model (ViT-B/16-RGB) [21], which is pre-trained on ImageNet.
We run the experiments with the model embeddings of size 768, binary embeddings, and two different hash encodings: The trivial hash function and Locality-Sensitive Hashing (LSH) [22]. LSH splits the latent space into two areas for each binary value using a random hyperplane. Both hash methods use a hash length of 32 to balance compression and accuracy. We do not vary the length in the reported results as other works already studied the effect [5, 6, 12]. Generally, decreasing the hash length reduces the precision.
@@ 4.2 Datasets
Reviewing remote sensing datasets revealed a limited availability of multi-spectral multi-class datasets with an image size of equal or more than 224 pixels [4, 23, 24]. BigEarthNet [17] and ForestNet [25] fulfill these requirements.
BigEarthNet consists of Sentinel-1 and -2 images with 120x120 pixels and includes two sets of multi-label annotations with 19 resp. 43 classes. The classes cover different land-use types such as mixed forest, water bodies, or airports. We are using the six Sentinel-2 channels supported by Prithvi. The images are bi-linear scaled up to match the model input, which corresponds to a five-meter resolution.
ForestNet includes Landsat 8 imagery from forest loss events. The images are annotated with twelve classes and four super-classes. The classes indicate types of deforestation, such as timber plantation or small-scale agriculture. ForestNet uses composite images created by averaging up to five cloud-free images. The images have a size of 332x332 and a 22-meter resolution after re-scaling.
Note that the data of both datasets differs from the Prithvi pre-training: The HLS data has a 30-meter resolution and includes additional data processing for harmonization. Furthermore, Prithvi-100M is pre-trained on US data and does not include any data from BigEarthNet and ForestNet which cover Europe and Indonesia, resp. [9, 17, 25].
Fig. 2: t-SNE plots of the ForestNet-4 test set with colored classes comparing the embedding space with hash codes.
Fig. 3: Examples from two datasets with query images (left), their labels, and retrieved images (right) using Prithvi and the trivial hash method. Images with green frames indicate positive matches, while those with red frames have different labels. Orange shows partial correct matches, where the number represents the number of label matches within the multi-labels.
@@ 5 Results
Our results are presented in Table 1. Prithvi-100M outperforms Prithvi-100M-RGB in every method by over 5.5 pp. on average. The differences for hash methods are significant for the BigEarthNet dataset, while the performance is comparable on ForestNet. The comparison of Prithvi-100M with the vanilla ViT-B/16-RGB leads to similar results, except a better-performing LSH method. LSH seems better suited for the vanilla ViT as it performs 5.61 pp. better than Prithvi-100M and on pair with the trivial hash method. Overall, the comparison highlights the information gain when using multi-spectral data with GeoFMs.
The transition from float embeddings to hash-based approaches typically results in a notable drop in model performance. E.g., Prithvi-100M with a 32-bit trivial hash has an 11.81 pp. lower mAP. This drop highlights the challenge of maintaining accuracy while simplifying the data representations. However, we observe that the performance loss is mainly influenced by the hash length and not by binarization. The binary embeddings perform nearly as well as the floating-point embeddings, with an average difference of only 0.62 pp. mAP for Prithvi.
We visually compare the ForestNet-4 embeddings and hash functions with t-SNE plots in Figure 2. The latent space is well distributed for the default and binarized embeddings but clustered with LSH. The clusters do not discriminate the semantic classes, which explains the lower performance. In comparison, trivial hash codes also include some larger clusters but retain the overall distribution better.
Figure 3 shows examples of retrieved images based on the trivial hashes. Retrieved images from BigEarthNet-43 have a large share of images with partial label matches, which is also represented by a 82.75 mAP. Those images are often similar in their overall color and appearance. However, some classes are less frequently matched, e.g., specific infrastructure classes like construction sites or airports. This is also observable in ForestNet-12, which has more fine-grained classes that are harder to differentiate. Still, the trivial approach returns some matching examples for most queries.
@@
Data type
Length
10K
50K
100K
Binary
32
16 ms
16 ms
17 ms
Binary
768
16 ms
16 ms
17 ms
Float
768
21 ms
32 ms
33 ms
@@
The retrieval speed for binary vectors is almost not influenced by the database size or the vector length. Floating-point embeddings have a retrieval time of up to 33 ms. This is two times longer than binary embeddings with 17 ms. The results demonstrate the efficient implementation of retrieval algorithms in vector databases like Milvus. Therefore, the hash length is mainly constrained by memory usage rather than the retrieval speed.
@@ 6 Conclusion
This work demonstrates the applicability of GeoFMs for image retrieval in remote sensing. Due to the learned representations through the pre-training, the model encodes multiple semantics and does not require further fine-tuning. We introduce two multi-spectral datasets to the retrieval task and provide strong baselines, enabling a more holistic evaluation for remote sensing in the future. Accordingly, we could evaluate three compression methods with binary embeddings, having shown the best trade-off between accuracy and retrieval speed. The approach can be easily implemented in various applications with any existing GeoFM and combined with more advanced compression methods to improve performance.

### v2

@@ 1 Introduction
Geospatial Foundation Models (GeoFM), for instance Prithvi [9] and SatMAE [10], open new possibilities because they are pre-trained on a vast amount of multi-spectral data. The models have been applied to various tasks, including flood and wildfire segmentation [9]. GeoFMs can utilize multi-spectral data for CBIR without requiring fine-tuning.
@@ 2 Related work
Image retrieval in remote sensing has evolved significantly, shifting focus from traditional metadata-based methods to more advanced techniques that use computer vision to analyze the image content [4]. Among these developments, deep hash networks have played a central role [5, 8, 11, 12]. These models compress images into hash codes which are stored in a database and used for calculating distances. Researchers explored various learning techniques, such as contrastive learning [8] and metric learning [13], with the aim of minimizing the need for extensive training data.
@@
GeoFMs, such as SatMAE, Prithvi, and Presto [19], have the ability to process multi-spectral data. Their pre-training on diverse satellite images overcomes the limitations of general models and benefits earth observation tasks [9].
@@ 3 Approach
The CBIR task requires identifying and retrieving images from a large database based on their visual and semantic similarity to a given query image. This process requires an efficient mechanism to compare and rank the database images in terms of their similarity to the query.
@@
Model Method BigEarthNet-43 BigEarthNet-19 ForestNet-12 ForestNet-4 Mean
@@
Binary emb. 97.44 97.83 43.28 59.85 74.60
64-bit hash 92.58 93.44 41.49 55.93 70.86
SatMAE-ViT-B Embedding 94.78 95.59 37.61 52.94 70.23
Binary emb. 89.39 90.40 36.49 53.04 67.33
64-bit hash 79.35 80.99 30.79 47.19 59.58
Prithvi-100M-RGB Embedding 92.15 93.17 38.65 53.85 69.46
Binary emb. 91.38 92.43 38.11 53.31 68.81
64-bit hash 82.60 84.45 32.58 48.20 61.96
ViT-B/16-RGB Embedding 89.31 90.21 38.92 56.49 68.73
Binary emb. 88.71 89.70 39.19 57.01 68.65
64-bit hash 79.01 81.54 33.60 49.63 60.95
Table 1: mAP@20 results for all models and datasets. We highlight the best-performing method in bold and underline the second-best one. The 64-bit hash uses the described trivial hash approach by averaging twelve embedding values for each bit.
@@ 4.1 Models
We evaluate Prithvi-100M and SatMAE-ViT-B as the underlying GeoFMs. Prithvi-100M is a Vision Transformer (ViT) [22] with 100M parameters and a 224x224 pixel input. Prithvi processes six bands, unlike the three RGB channels of the vanilla ViT. The pre-training includes a subset of the Harmonized Landsat-Sentinel (HLS) dataset consisting of images from the Landsat 8 and Sentinel-2 satellites at 30 meter resolution. We also report the results of Prithvi only processing the RGB channels for comparison (Prithvi-100M-RGB). The infrared channels are set to the channel mean which is mathematically similar to dropping these channels. SatMAE is a ViT-B/16 model with a 96x96 pixel input. The model is trained on ten Sentinel-2 bands at 10 meter resolution. For Landsat images, we fill the missing bands with the channel mean similar to Prithvi-100M-RGB. For comparison, we include the vanilla ViT-B/16 model (ViT-B/16-RGB) [22], which is pre-trained on ImageNet-21K.
We run the experiments with the model embeddings of size 768, binary embeddings, and trivial hash codes of length 64. We varied the hash length in additional experiments and find that 64 bit is a good compromise between compression and accuracy, as the performance drops significantly with a smaller hash length. We also tested Locality-Sensitive Hashing (LSH) [23] but did not observed a overall improved performance compared to the trivial approach.
@@ 4.2 Datasets
Reviewing remote sensing datasets revealed a limited availability of multi-spectral multi-class datasets with an sufficient image size [4, 24, 25]. BigEarthNet [18] and ForestNet [26] fulfill these requirements.
BigEarthNet consists of Sentinel-1 and Sentinel-2 images with 120x120 pixels and includes two sets of multi-label annotations with 19 resp. 43 classes. The classes cover different land-use types such as mixed forest, water bodies, or airports. We are using the six resp. ten Sentinel-2 channels supported by Prithvi and SatMAE.
ForestNet includes Landsat 8 imagery from forest loss events with a 332x332 image size. The images are annotated with twelve classes and four super-classes. The classes indicate types of deforestation, such as timber plantation or small-scale agriculture. ForestNet uses composite images created by averaging up to five cloud-free images.
Note that the data of both datasets differs from the pre-training settings. The images are bi-linear re-scaled to match the model input size, which leads to different spatial resolutions than during pre-training. Furthermore, SatMAE was not trained on Landsat data and Prithvi-100M is pre-trained only on US data while BigEarthNet and ForestNet cover Europe and Indonesia, resp. [9, 18, 26].
Fig. 2: t-SNE plots of the ForestNet-4 test set with colored classes comparing Pritvi-100M embeddings.
Fig. 3: Examples from two datasets with query images (left), their labels, and retrieved images (right) using Prithvi-100M and the 64-bit trivial hash. Images with green frames indicate positive matches, while those with red frames have different labels. Orange shows partial correct matches, where the number represents the number of label matches within the multi-labels.
@@ 5 Results
Prithvi-100M outperforms all other models in every method by over 4.9 pp. on average as shown in Table 1. The transition from float embeddings to hash-based approaches results in a notable drop in model performance. E.g., Prithvi-100M with a 64-bit trivial hash has on average an 4.36 pp. lower mAP. The difference is larger for other models: SatMAE and vanilla ViT have a average difference of 10.65 and 7.78 pp. resp. This drop highlights the challenge of maintaining accuracy while simplifying the data representations. Except for SatMAE, we observe that the performance loss is mainly influenced by the hash length and not by binarization. The binary embeddings perform nearly as well as the floating-point embeddings, with an average difference of only 0.62 pp. mAP for Prithvi-100M.
While SatMAE outperforms the RGB-only models on BigEarthNet, it performs worse on ForestNet. The missing channels and different data source might affect the performance. Overall, the comparison highlights the information gain when using multi-spectral data with GeoFMs that are trained on satellite data.
We visually compare the ForestNet-4 embeddings and hash functions with t-SNE plots in Figure 2. The latent space is well distributed for the float and binarized embeddings as well as the 64-bit trivial hash. The 32-bit trivial hash has clusters which do not discriminate the semantic classes and explain the much lower performance of shorter hash codes.
Figure 3 shows examples of retrieved images based on the trivial hashes. Retrieved images from BigEarthNet-43 have a large share of images with partial label matches, which is also represented by a 92.58 mAP. Those images are often similar in their overall color and appearance. However, some classes are rarely matched, e.g., specific infrastructure classes like construction sites or airports. This is also observable in ForestNet-12, which has more fine-grained classes that are harder to differentiate. Still, our approach returns some matching examples for most queries from this dataset.
@@
Data type Length 10K 50K 100K
Binary 64 16 ms 16 ms 17 ms
Binary 768 16 ms 16 ms 17 ms
Float 768 21 ms 32 ms 33 ms
@@
The retrieval speed for binary vectors is almost not influenced by the database size or the vector length. Floating-point embeddings have a retrieval time of up to 33 ms. This is two times longer than binary embeddings with 17 ms. The results demonstrate the efficient implementation of retrieval algorithms in vector databases like Milvus by using cluster centers. Therefore, the hash length is mainly constrained by memory usage rather than the retrieval speed.
@@ 6 Conclusion
This work demonstrates the applicability of GeoFMs for image retrieval in remote sensing. Due to the learned representations through the pre-training, Prithvi-100M encodes multiple semantics and does not require further fine-tuning. We introduce two multi-spectral datasets to the retrieval task and provide strong baselines, enabling a more holistic evaluation for remote sensing in the future. Accordingly, we evaluated two compression methods and binary embeddings have shown the best trade-off between accuracy and retrieval speed. The approach can be easily implemented in various applications with any existing GeoFM and combined with more advanced compression methods to improve performance.